OpenStudy (mackenzie6302):

Write 5.2 as the quotient of two integers. 11/2 51/10 26/5 21/5

OpenStudy (welshfella):

5.2 = 5 1/5

OpenStudy (welshfella):

5 1/5 = what /

OpenStudy (mathstudent55):

5.2 = 52/10 Now reduce the fraction.

OpenStudy (isaidavila):

5.2= 52/10=?

OpenStudy (jdoe0001):

and so, what did you get?